OpenStudy (anonymous):

Could someone help me please? 3x-

OpenStudy (anonymous):

3s minus what?

OpenStudy (anonymous):

3x - 4y =5, 2x - y = 1

OpenStudy (anonymous):

To sole these problems, we find one variable (say x) in terms of the other in one equation, then chuck it into the other equation. First, lets take: \[2x - y =1 \] and make it in terms of y. So we have: \[ y = 2x-1\] Now in the first equation: \[3x - 4y = 5\] We substitute y: \[3x -4(2x -1)=5\] Upon solving this we have: \[3x -8x+4=5\] \[-5x +4 = 5\] \[-5x=1\] \[x=-\frac{1}{5}\] Now we have the value of x. Can you find the value of y yourself?

OpenStudy (anonymous):

You way also want to look at: www.youtube.com/watch?v=cwHR_B9zK7k

OpenStudy (anonymous):

Could you maybe get me started, I've been out of school for 24 years and I'm really trying to understand this new generation math

OpenStudy (anonymous):

Take a look at the video, I find patrickJMT videos to be really good.

OpenStudy (anonymous):

Okay,thank you so much for your help

OpenStudy (anonymous):

If you still need help, I'll be back in half an hour. Good luck!

OpenStudy (anonymous):

Thanks